OpenStudy (amber15):

A rectangular table top measures 5/8 of a meter long by 1/7 of a meter wide . what is the area ? The area is ___ square meters

OpenStudy (anonymous):

area is length * width. simply multiply through: \[\frac{ 5 }{ 8 } * \frac{ 1 }{ 7 } =\frac{ 5 }{ 56 }\]

OpenStudy (anonymous):

.089 m^2

OpenStudy (amber15):

oh okay im always lost on length width and height not good in math